if a block is 6cm in length 5cm in breadth and 2cm in height and you double the sides as before do you find a similar relationship between the volume of the block and the volume of the block with doubled sides?
Answer by rajagopalan(148)

Block Size=6*5*2=60 Cu cm
New block=12*10*4=480 Cu cm
Vol increas ratio = 480/50=8 times
When sides are doubled the volume increases by 8 times
The relationship is not similar but rises by 8 times (2^3).